OpenStudy (anonymous):

Find the equation of the circle with center at (-3, 1) and through the point (2, 13)

OpenStudy (jdoe0001):

once you have the center, you have the (h,k) values for your circle, then all you'd need is the radius

OpenStudy (anonymous):

How would i find the radius?

OpenStudy (jdoe0001):

http://www.ipracticemath.com/Content/imagesLM/Algebra/distance-formula1.png the distance formula, between 2 points, you do have 2 points here, will give you the diameter, half that is the radius

OpenStudy (anonymous):

So would the final equation be (x+3)2+(y-1)2=121 ?

OpenStudy (jdoe0001):

yes, but I don't get 121 though

OpenStudy (jdoe0001):

lemme recheck

OpenStudy (jdoe0001):

(-3-2) = 5 (1-13) = 12

OpenStudy (jdoe0001):

so, that makes it \(\sqrt{25+144} \)

OpenStudy (anonymous):

Ok i see where i went wrong, that makes it so the diameter would be 13 and the radius would be 6.5

OpenStudy (jdoe0001):

(-3-2) = -5, but anyhow, (-5)^2 still is 25

OpenStudy (jdoe0001):

so,the distance between those 2 points is "13", and the radius would be 13/2 so \( (x+3)^2+(y-1)^2=\pmatrix{\frac{13}{2}}^2 \)

OpenStudy (anonymous):

okay thank you! now i understand it way better :)

OpenStudy (jdoe0001):

yw